OpenStudy (anonymous):

A,B,C,D start at the same time to run around a circular garden in the same direction .A completes a round in 30 minutes,B in 60 minutes,c in 90 minutes and D in 105 minutes.after what time will they meet again at the starting point

OpenStudy (anonymous):

you should find lcm(30, 60, 90, 105)

OpenStudy (anonymous):

is it 15

OpenStudy (anonymous):

if i tell i will be banned

OpenStudy (anonymous):

but it isn't

OpenStudy (anonymous):

k.5,3,3,4,3,7 is the lcm ., as multiply=3280

OpenStudy (debbieg):

@pooja123 , when you say 15 you are thinking of GCF (greatest common factor). You want LCM - LCM is ALWAYS at least as LARGE as the largest number.

OpenStudy (anonymous):

nope, dude you should take 3 only one time

OpenStudy (anonymous):

k.5,3,7 ,4=420min

OpenStudy (anonymous):

420 divided by 90 =?

OpenStudy (debbieg):

There is a brute force way to find LCM, and an algorithmic way. Brute force: list multiples of each number, until you find the smallest number that is common to all. Can be tedious with several numbers. Algorithmic way: do the prime factorization of each number. Then "cook up" the GCF as the product of ALL the factors, the MOST number of times they appear in any of the factorizations.

OpenStudy (debbieg):

E.g., if I want the LCM of 40,50 I do the prime factors of each: 40=2*2*2*5 50=2*5*5 Then my LCM need to have three factors of 2, and two factors of 5: LCM=2*2*2*5*5=200

OpenStudy (debbieg):

Now apply that to your four numbers above. It is guaranteed to work every time. :)

OpenStudy (anonymous):

k.i got it.

OpenStudy (debbieg):

okeydoke, good.

OpenStudy (anonymous):

2,3,5,7=210

OpenStudy (anonymous):

sry 3^2,2^2,5,7=1860

OpenStudy (debbieg):

no. you have an extra factor of 2 there, and your product isn't right regardless (e.g., you don't get 1860 from that product).

OpenStudy (debbieg):

You only need 1 factor of 2, because none of the numbers have 4 as a factor.

OpenStudy (debbieg):

Did you do the prime factorizations? (And really, since 90 is a multiple of 30 and 60, you ONLY need to find GCF(90,105), the factors of 30 and 60 are all included in the factors of 90)

OpenStudy (anonymous):

210

OpenStudy (debbieg):

So really, you just need 90 * (any factors of 105 that are NOT factors of 90)

OpenStudy (debbieg):

No, too small. 90 is not a factor of 210.

OpenStudy (debbieg):

If you show me your work, I can help you understand why it isn't working. If you follow the algorithm I've explained, it will work.

OpenStudy (debbieg):

Give the prime factorizations of: 90: 105:

OpenStudy (anonymous):

30=3.2.5 60=3.2.25 90=3.2.3.5 105=3.7.5

OpenStudy (debbieg):

Oh crap... you're right, I was forgetting about the 2*2 in 60. lol duh to me. So it does not reduce to LCM(90,105) as I said, but it does reduce to LCM(60,90,105) - you can basically ignore the 30 because any multiple of 60 or 90 is also a multiple of 30.

OpenStudy (anonymous):

1260

OpenStudy (debbieg):

Right.... your set up above 3^2,2^2,5,7=1860 is right, just the result was wrong. Sorry about my "4" snafu! But you have the right idea now. : )

OpenStudy (anonymous):

k.thanks a lot

OpenStudy (debbieg):

you're welcome. :)